OpenStudy (anonymous):

lim x->-infinite 1/x = 0 Could someone explain why it is equal to zero?

OpenStudy (zzr0ck3r):

Is this for regular calculus or analysis?

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} 1/x=0\]

OpenStudy (zzr0ck3r):

1/10=.1 1/100=.01 1/1000000000000000000000 =.000000000000000000001

OpenStudy (anonymous):

Its just for calc, I don't understand why its zero

OpenStudy (anonymous):

oh is that why?

OpenStudy (zzr0ck3r):

yep

OpenStudy (anonymous):

could you show me on a graph?

OpenStudy (zzr0ck3r):

you can get as close as you want to 0, by putting a bigger number on the bottom

OpenStudy (zzr0ck3r):



OpenStudy (zzr0ck3r):

the larger x gets, the closer it gets to 0

OpenStudy (anonymous):

oh ok I see. what would it be if it was x^2

OpenStudy (zzr0ck3r):

what happens when you put larger numbers in x^2?

OpenStudy (anonymous):

they double

OpenStudy (loser66):

it's not double, it's exponent

OpenStudy (anonymous):

so it would be infinite

OpenStudy (zzr0ck3r):

correct